# The LSCD Benchmark: a Testbed for Diachronic Word Meaning Tasks

### v1

@@ Abstract
Lexical Semantic Change Detection (LSCD) is a complex, lemma-level task, which is usually operationalized based on two subsequently applied usage-level tasks: First, Word-in-Context (WiC) labels are derived for pairs of usages. Then, these labels are represented in a graph on which Word Sense Induction (WSI) is applied to derive sense clusters. Finally, LSCD labels are derived by comparing sense clusters over time. This modularity is reflected in most LSCD datasets and models. It also leads to a large heterogeneity in modeling options and task definitions, which is exacerbated by a variety of dataset versions, preprocessing options and evaluation metrics. This heterogeneity makes it difficult to evaluate models under comparable conditions, to choose optimal model combinations or to reproduce results. Hence, we provide a benchmark repository standardizing LSCD evaluation. Through transparent implementation results become easily reproducible and by standardization different components can be freely combined. The repository reflects the task’s modularity by allowing model evaluation for WiC, WSI and LSCD. This allows for careful evaluation of increasingly complex model components providing new ways of model optimization.
1 Introduction
@@
In order to handle this heterogeneity, we think that a shared testbed with a common evaluation setup is needed. Hence, we present a benchmark repository implementing evaluation procedures for models on most available LSCD datasets. [^1] https://github.com/ChangeIsKey/LSCDBenchmark The benchmark exploits the modularity of the meta task LSCD by allowing for evaluation of the subtasks WiC and WSI on the same datasets. It can be assumed that performance on the subtasks directly determines performance on the meta task. We aim to stimulate transfer between the fields of WiC, WSI and LSCD by providing a repository allowing for evaluation on all these tasks with shared model components.
@@
We hope that the resulting benchmark by standardizing the evaluation of LSCD models and providing models with near-SOTA performance can serve as a starting point for researchers to develop and improve models. The benchmark allows for a wide application and testing of models by focusing on multilingual models and their evaluation on several languages.
2 Related Work
@@
A number of recently created LSCD data sets apply WiC and WSI in the annotation process (cf. Section 3) and thus allow for evaluation of WiC and WSI along with LSCD models (i.a. Schlechtweg et al., 2021; Kurtyigit et al., 2021; Kutuzov et al., 2022; Zamora-Reina et al., 2022). [^2] The bulk of these data sets is listed at: https://www.ims.uni-stuttgart.de/data/wugs. There are also a number of data sets omitting WSI, but allowing for WiC and LSCD evaluation (i.a. Schlechtweg et al., 2018; Rodina and Kutuzov, 2020; Kutuzov and Pivovarova, 2021), or data sets omitting the WiC allowing for WSI and LSCD evaluation (i.a. Basile et al., 2020; Cook et al., 2014), or datasets allowing purely for WiC evaluation (Loureiro et al., 2022). Above that, there is a number of data sets allowing for synchronic evaluation of WiC (i.a. Pilehvar and Camacho-Collados, 2019; Trott and Bergen, 2021) or WSI (i.a. Langone et al., 2004; Hovy et al., 2006) or both (Erk et al., 2013; Aksenova et al., 2022), which can be exploited e.g. by simulating LSCD labels (Rosenfeld and Erk, 2018; Dubossarsky et al., 2019; Schlechtweg and Schulte im Walde, 2020). Note also that there is no restriction on the source or strategy for sampling word usages, i.e., they do not necessarily have to be sampled from different time periods, but could also be sampled from different text genres, domains, dialects or even languages. Hence, data sets reflecting meaning divergences between these text categories can be integrated into the benchmark as well (i.a. Hätty et al., 2019; Baldissin et al., 2022).
@@
So far, there is no comprehensive LSCD benchmark, implementing state-of-the-art models on (human-annotated) high-quality evaluation data from multiple languages and multiple time periods. The leaderboards of several standalone shared tasks can be seen as small-scale benchmarks without common model implementation (Ahmad et al., 2020; Basile et al., 2020; Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022) with the SemEval task being the most diverse with four languages (Schlechtweg et al., 2020). Schlechtweg et al. (2019) provide a comprehensive repository of type-based modeling approaches to LSCD with evaluation pipelines on multiple data sets. [^3] https://github.com/Garrafao/LSCDetection However, type-based models have more recently been outperformed by token-based contextualized embedding approaches (Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022; Cassotti et al., 2023). Moreover, Duong et al. (2021) provide a repository to generate synthetic evaluation data for the related task of discourse shift detection with several model implementations not representing the state-of-the-art in LSCD. [^4] https://github.com/ruathudo/detangling-discourses
@@ 3 Tasks
3.1 Word-in-Context
The Word-in-Context task is to determine if two words occurring in two text fragments have the same or different meanings. Usually two usages of the same word probably in different grammatical forms are given. [^7] However, there are datasets with examples consisting of usages of two different words that are similar in one of their meanings (Huang et al., 2012; Armendariz et al., 2020; Baldissin et al., 2022), and in the cross-lingual setup these two words and the corresponding text fragments are in different languages (Martelli et al., 2021). For example: {example} Von Hassel replied that he had such faith in the plane that he had no hesitation about allowing his only son to become a Starfighter pilot. {example} This point, where the rays pass through the perspective plane, is called the seat of their representation.
The WiC task is often framed as a binary classification task. For instance, the WiC (Pilehvar and Camacho-Collados, 2019) and MCL-WiC (Martelli et al., 2021) datasets contain binary labels and employ accuracy as the main evaluation metric. Alternatively, USim (Erk et al., 2013), SCWS (Huang et al., 2012) and CoSimLex (Armendariz et al., 2020) were labeled with non-binary similarity scores and promote a graded formulation of the task. In this formulation, a WiC model shall produce scores that are similar to the human scores, or at least rank the pairs of usages similarly. Spearman’s and Pearson’s correlation coefficients are employed as evaluation metrics in this case.
Following the DURel annotation framework (Schlechtweg et al., 2024), during the annotation process of most LSCD datasets human annotators were essentially solving the graded WiC task, i.e. they annotated the similarity of two usages of the same word on a scale. This provides data for evaluation of WiC models that can serve as a part of LSCD models. In diachronic LSCD datasets there are pairs of word usages extracted from two documents belonging to distant time periods making usages in these pairs very different orthographically, grammatically, and thematically, even when the target word has the same meaning. This might be challenging for models trained on traditional WiC datasets, which often contain examples from the same time period. Our benchmark helps to analyze how sensitive WiC models are to this shift in time period by comparing their performance on pairs of usages extracted from the old, the new or both corpora.
3.2 Word Sense Induction
@@
Data set LGS $\mathbf{n}$ N/V/A $\mathbf{|U|}$ AN JUD Task $\mathbf{t_{1}}$ $\mathbf{t_{2}}$ Reference Version DWUG DE 48 32/14/2 178 8 37k WiC, WSI, LSCD (B,G,C) 1800–1899 1946–1990 Schlechtweg et al. (2021) 2.2.0 DWUG EN 40 36/4/0 189 9 29k WiC, WSI, LSCD (B,G,C) 1810–1860 1960–2010 Schlechtweg et al. (2021) 2.0.1 DWUG SV 40 31/6/3 168 5 20k WiC, WSI, LSCD (B,G,C) 1790–1830 1895–1903 Schlechtweg et al. (2021) 2.0.1 DWUG ES 100 51/24/25 40 12 62k WiC, WSI, LSCD (B,G,C) 1810–1906 1994–2020 Zamora-Reina et al. (2022) 4.0.0 DiscoWUG DE 75 39/16/20 49 8 24k WiC, WSI, LSCD (B,G,C) 1800–1899 1946–1990 Kurtyigit et al. (2021) 1.1.1 RefWUG DE 22 15/1/6 19 5 4k WiC, WSI, LSCD (B,G,C) 1750–1800 1850–1900 Schlechtweg (2023) 1.1.0 NorDiaChange1 NO 40 40/0/0 21 3 14k WiC, WSI, LSCD (B,G,C) 1929–1965 1970–2013 Kutuzov et al. (2022) 1.0.0 NorDiaChange2 NO 40 40/0/0 21 3 15k WiC, WSI, LSCD (B,G,C) 1980–1990 2012–2019 Kutuzov et al. (2022) 1.0.0 DURel DE 22 15/1/6 104 5 6k WiC, LSCD (C) 1750–1800 1850–1900 Schlechtweg et al. (2018) 3.0.0 SURel DE 22 19/3/0 104 4 5k WiC, LSCD (C) general domain Hätty et al. (2019) 3.0.0 RuSemShift1 RU 71 65/6/0 119 5 21k WiC, LSCD (C) 1682–1916 1918–1990 Rodina and Kutuzov (2020) 2.0.0 RuSemShift2 RU 69 57/12/0 105 5 18k WiC, LSCD (C) 1918–1990 1991–2016 Rodina and Kutuzov (2020) 2.0.0 RuShiftEval1 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1682–1916 1918–1990 Kutuzov and Pivovarova (2021) 2.0.0 RuShiftEval2 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1918–1990 1991–2016 Kutuzov and Pivovarova (2021) 2.0.0 RuShiftEval3 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1682–1916 1991–2016 Kutuzov and Pivovarova (2021) 2.0.0
@@
3.3 Lexical Semantic Change Detection
@@
The JSD task has the same assumptions about word meaning, but instead of binary classification it requires ranking a given list of words according to changes in their sense frequency distributions. The rank of a word is determined by the Jensen–Shannon Distance between two probability distributions over word senses $P(sense|w,t_{old})$ and $P(sense|w,t_{new})$, one for the older time period and another for the newer one (Schlechtweg et al., 2020; Zamora-Reina et al., 2022).
Finally, the COMPARE task requires ranking a given set of words according to the average similarity by meaning between old and new usages of each word (Schlechtweg et al., 2018; Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022). The COMPARE task can be reduced to the graded WiC task for the pairs of usages consisting of one old and one new usage, the final word scores can be obtained by averaging WiC scores of these pairs.
@@ 5 Evaluation Procedures
The central part shows the WSI-based approach to LSCD. It relies on WSI methods which cluster word usages based either on their contextualized embeddings or pairwise similarities between them calculated by a WiC model. If a WiC model is involved, we can evaluate it separately on all datasets containing human-labeled pairs of word usages by feeding these pairs and comparing the model predictions with the human labels. Spearman’s and Person’s correlation coefficients that compare rankings or scores predicted by humans and the WiC model are employed as metrics for the WiC task.
@@
6 Conclusion

### v2

@@ Abstract
Lexical Semantic Change Detection (LSCD) is a complex, lemma-level task, which is usually operationalized based on two subsequently applied usage-level tasks: First, Word-in-Context (WiC) labels are derived for pairs of usages. Then, these labels are represented in a graph on which Word Sense Induction (WSI) is applied to derive sense clusters. Finally, LSCD labels are derived by comparing sense clusters over time. This modularity is reflected in most LSCD datasets and models. It also leads to a large heterogeneity in modeling options and task definitions, which is exacerbated by a variety of dataset versions, preprocessing options and evaluation metrics. This heterogeneity makes it difficult to evaluate models under comparable conditions, to choose optimal model combinations or to reproduce results. Hence, we provide a benchmark repository standardizing LSCD evaluation. Through transparent implementation results become easily reproducible and by standardization different components can be freely combined. The repository reflects the task’s modularity by allowing model evaluation for WiC, WSI and LSCD. This allows for careful evaluation of increasingly complex model components providing new ways of model optimization. We use the implemented benchmark to conduct a number of experiments with recent models and systematically improve the state-of-the-art.
Keywords: Lexical Semantic Change Detection, Word-in-Context, Word Sense Induction, Word Usage Graphs, Diachronic, Word Meaning
[…]
The LSCD Benchmark: a Testbed for Diachronic Word Meaning Tasks
Dominik Schlechtweg1, Sachin Yadav1, Nikolay Arefyev2
1University of Stuttgart, 2University of Oslo
first.last@ims.uni-stuttgart.de, nikolare@uio.no
Abstract content
1.  Introduction
@@
In order to handle this heterogeneity, we think that a shared testbed with a common evaluation setup is needed. Hence, we present a benchmark repository implementing evaluation procedures for models on most available LSCD datasets. [^1] Find the code at https://github.com/Garrafao/LSCDBenchmark. The benchmark exploits the modularity of the meta task LSCD by allowing for evaluation of the subtasks WiC and WSI on the same datasets. It can be assumed that performance on the subtasks directly determines performance on the meta task. We aim to stimulate transfer between the fields of WiC, WSI and LSCD by providing a repository allowing for evaluation on all these tasks with shared model components.
@@
We hope that the resulting benchmark by standardizing the evaluation of LSCD models and providing models with SOTA performance can serve as a starting point for researchers to develop and improve models. The benchmark allows for a wide application and testing of models by focusing on multilingual models and their evaluation on several languages.
2.  Related Work
@@
A number of recently created LSCD datasets apply WiC and WSI in the annotation process (cf. Section 3) and thus allow for evaluation of WiC and WSI along with LSCD models (i.a. Schlechtweg et al., 2021; Kurtyigit et al., 2021; Kutuzov et al., 2022; Zamora-Reina et al., 2022; Chen et al., 2023). [^2] The bulk of these datasets is listed at: https://www.ims.uni-stuttgart.de/data/wugs. There are also a number of datasets omitting WSI, but allowing for WiC and LSCD evaluation (i.a. Schlechtweg et al., 2018; Rodina and Kutuzov, 2020; Kutuzov and Pivovarova, 2021), or datasets omitting the WiC allowing for WSI and LSCD evaluation (i.a. Basile et al., 2020; Cook et al., 2014), or datasets allowing purely for WiC evaluation (Loureiro et al., 2022). Above that, there is a number of datasets allowing for synchronic evaluation of WiC (i.a. Pilehvar and Camacho-Collados, 2019; Trott and Bergen, 2021) or WSI (i.a. Langone et al., 2004; Hovy et al., 2006) or both (Erk et al., 2013; Aksenova et al., 2022), which can be exploited e.g. by simulating LSCD labels (Rosenfeld and Erk, 2018; Dubossarsky et al., 2019; Schlechtweg and Schulte im Walde, 2020). Note also that there is no restriction on the source or strategy for sampling word usages, i.e., they do not necessarily have to be sampled from different time periods, but could also be sampled from different text genres, domains, dialects or even languages. Hence, datasets reflecting meaning divergences between these text categories can be integrated into the benchmark as well (i.a. Hätty et al., 2019; Baldissin et al., 2022).
@@
So far, there is no comprehensive LSCD benchmark, implementing state-of-the-art models on (human-annotated) high-quality evaluation data from multiple languages and multiple time periods. The leaderboards of several standalone shared tasks can be seen as small-scale benchmarks without common model implementation (Ahmad et al., 2020; Basile et al., 2020; Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022; Fedorova et al., 2024) with the SemEval task being the most diverse with four languages (Schlechtweg et al., 2020). Schlechtweg et al. (2019) provide a comprehensive repository of type-based modeling approaches to LSCD with evaluation pipelines on multiple datasets. [^3] https://github.com/Garrafao/LSCDetection However, type-based models have more recently been outperformed by token-based contextualized embedding approaches (Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022; Cassotti et al., 2023). Periti and Tahmasebi (2024) perform a systematic comparison, but do not provide a flexible model implementation and do not evaluate on some of the most recent SOTA models or datasets. Moreover, Duong et al. (2021) provide a repository to generate synthetic evaluation data for the related task of discourse shift detection with several model implementations not representing the state-of-the-art in LSCD. [^4] https://github.com/ruathudo/detangling-discourses
@@ 3.  Tasks
3.1.  Word-in-Context
The Word-in-Context task is to determine if two words occurring in two text fragments have the same or different meanings. Usually two usages of the same word probably in different grammatical forms are given. [^7] However, there are datasets with examples consisting of usages of two different words that are similar in one of their meanings (Huang et al., 2012; Armendariz et al., 2020; Baldissin et al., 2022), and in the cross-lingual setup these two words and the corresponding text fragments are in different languages (Martelli et al., 2021). For example: {example} Von Hassel replied that he had such faith in the plane that he had no hesitation about allowing his only son to become a Starfighter pilot. {example} This point, where the rays pass through the perspective plane, is called the seat of their representation. The WiC task is often framed as a binary classification task. For instance, the WiC (Pilehvar and Camacho-Collados, 2019) and MCL-WiC (Martelli et al., 2021) datasets contain binary labels and employ accuracy as the main evaluation metric. Alternatively, USim (Erk et al., 2013), SCWS (Huang et al., 2012) and CoSimLex (Armendariz et al., 2020) were labeled with non-binary semantic proximity scores and promote a graded formulation of the task. In this formulation, a WiC model shall produce scores that are similar to the human scores, or at least rank the pairs of usages similarly. Spearman’s and Pearson’s correlation coefficients are employed as evaluation metrics in this case. More recently, Schlechtweg et al. (2025) provided an ordinal formulation of the graded WiC task evaluating with Krippendorff’s $\alpha$ (Krippendorff, 2018).
Following the DURel annotation framework (Schlechtweg et al., 2024b), during the annotation process of most LSCD datasets human annotators were essentially solving the graded WiC task, i.e., they annotated the semantic proximity of two usages of the same word on a scale. This provides data for evaluation of WiC models that can serve as a part of LSCD models. In diachronic LSCD datasets, there are pairs of word usages extracted from two documents belonging to distant time periods making usages in these pairs very different orthographically, grammatically, and thematically, even when the target word has the same meaning. This might be challenging for models trained on traditional WiC datasets, which often contain examples from the same time period. Our benchmark helps to analyze how sensitive WiC models are to this shift in time period by comparing their performance on pairs of usages extracted from the old, the new or both corpora.
3.2.  Word Sense Induction
@@
Data set LGS $\mathbf{n}$ N/V/A $\mathbf{|U|}$ AN JUD Task $\mathbf{t_{1}}$ $\mathbf{t_{2}}$ Reference Version DWUG DE 50 32/14/2 178 8 63k WiC, WSI, LSCD (B,G,C) 1800–1899 1946–1990 Schlechtweg et al. (2021) 3.0.0 DWUG Res. DE 15 10/4/1 50 3 10k WiC, WSI, LSCD (B,G,C) 1800–1899 1946–1990 Schlechtweg et al. (2024a) 1.0.0 DWUG EN 46 40/6/0 191 13 29k WiC, WSI, LSCD (B,G,C) 1810–1860 1960–2010 Schlechtweg et al. (2021) 3.0.0 DWUG Res. EN 15 14/1/0 50 3 7K WiC, WSI, LSCD (B,G,C) 1810–1860 1960–2010 Schlechtweg et al. (2024a) 1.0.0 DWUG SV 44 32/5/7 171 13 20k WiC, WSI, LSCD (B,G,C) 1790–1830 1895–1903 Schlechtweg et al. (2021) 3.0.0 DWUG Res. SV 15 10/3/2 50 6 16k WiC, WSI, LSCD (B,G,C) 1790–1830 1895–1903 Schlechtweg et al. (2024a) 1.0.0 DWUG ES 100 51/24/25 40 12 62k WiC, WSI, LSCD (B,G,C) 1810–1906 1994–2020 Zamora-Reina et al. (2022) 4.0.2 DiscoWUG DE 75 39/16/20 49 8 24k WiC, WSI, LSCD (B,G,C) 1800–1899 1946–1990 Kurtyigit et al. (2021) 2.0.0 RefWUG DE 22 15/1/6 19 5 4k WiC, WSI, LSCD (B,G,C) 1750–1800 1850–1900 Schlechtweg (2023) 1.1.0 NorDiaChange1 NO 40 40/0/0 21 3 14k WiC, WSI, LSCD (B,G,C) 1929–1965 1970–2013 Kutuzov et al. (2022) 1.0.0 NorDiaChange2 NO 40 40/0/0 21 3 15k WiC, WSI, LSCD (B,G,C) 1980–1990 2012–2019 Kutuzov et al. (2022) 1.0.0 ChiWUG ZH 40 10/22/8 40 4 61k WiC, WSI, LSCD (B,G,C) 954-1978 1979-2003 Chen et al. (2023) 1.0.0 DWUG IT 5k WiC, WSI, LSCD (B,G,C) 1948-1970 1990-2014 Cassotti et al. (2024) 3.0.0 DURel DE 22 15/1/6 104 5 6k WiC, LSCD (C) 1750–1800 1850–1900 Schlechtweg et al. (2018) 3.0.0 SURel DE 22 19/3/0 104 4 5k WiC, LSCD (C) general domain Hätty et al. (2019) 3.0.0 RuSemShift1 RU 71 65/6/0 119 5 21k WiC, LSCD (C) 1682–1916 1918–1990 Rodina and Kutuzov (2020) 2.0.0 RuSemShift2 RU 69 57/12/0 105 5 18k WiC, LSCD (C) 1918–1990 1991–2016 Rodina and Kutuzov (2020) 2.0.0 RuShiftEval1 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1682–1916 1918–1990 Kutuzov and Pivovarova (2021) 2.0.0 RuShiftEval2 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1918–1990 1991–2016 Kutuzov and Pivovarova (2021) 2.0.0 RuShiftEval3 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1682–1916 1991–2016 Kutuzov and Pivovarova (2021) 2.0.0
@@
$3.$  Lexical Semantic Change Detection
@@
The JSD (or Graded Change) task has the same assumptions about word meaning, but instead of binary classification it requires ranking a given list of words according to changes in their sense frequency distributions. The rank of a word is determined by the Jensen–Shannon Distance between two probability distributions over word senses $P(sense|w,t_{old})$ and $P(sense|w,t_{new})$, one for the older time period and another for the newer one (Schlechtweg et al., 2020; Zamora-Reina et al., 2022).
Finally, the COMPARE task requires ranking a given set of words according to the average proximity by meaning between old and new usages of each word (Schlechtweg et al., 2018; Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022). The COMPARE task can be reduced to the graded WiC task for the pairs of usages consisting of one old and one new usage, the final word scores can be obtained by averaging WiC scores of these pairs.
@@ 5.  Evaluation Procedures
The central part shows the WSI-based approach to LSCD. It relies on WSI methods which cluster word usages based either on their contextualized embeddings or pairwise similarities between them calculated by a WiC model. If a WiC model is involved, we can evaluate it separately on all datasets containing human-labeled pairs of word usages by feeding these pairs and comparing the model predictions with the human labels. Spearman’s and Pearson’s correlation coefficients that compare rankings or scores predicted by humans and the WiC model are employed as metrics for the WiC task.
@@
We introduce standard splits for each dataset on the lemma level, i.e., certain target words are assigned to train/dev/test. We further provide possibility to evaluate on the previously introduced standard split from the CoMeDi shared task (Schlechtweg et al., 2025).
6.  Models
The most straightforward modeling approach for LSCD models is to follow the 3-level annotation approach of WUGs described in Section 3. Given a target word, a basic model retrieves uses of this word from an old and a new corpus, then clusters them in order to infer word senses, and finally analyzes the obtained clusters to make conclusions about changes in word senses between two time periods. This approach is appealing because if word senses are inferred correctly, then an exact description of how word senses changed, as well as the exact predictions for all LSCD tasks are easy to obtain. Also the procedure basically automates the annotation procedure of various LSCD datasets, which is a reasonable way of getting predictions that correspond to the ground truth well. The benchmark implementation as depicted in Figure 2 thus follows this basic structure. The inputs and the outputs of each component are specified in round brackets, while the hyperparameters are underlined in the corresponding descriptions. The implemented components on each of these levels will be described below. It is important to note that current SOTA models for graded change (Giulianelli et al., 2020; Arefyev et al., 2021a; Cassotti et al., 2023) skip the clustering step and model graded change directly from Word-in-Context predictions of proximity between word usages or their underlying vectors by aggregating these time-wise (see below).
Figure 3: Result overview on Graded Change (left) and COMPARE score (right). XLD = XL-DURel, XLM = XL-LEXEME, MCL/MCLen = DeepMistake checkpoints, TH = thresholded, APD = Average Pairwise Distance, DIA = DiaSense, COS = Cosine distance between average embeddings.
Retrieve uses
(word, corpus → uses). Given a corpus and a target word, this component retrieves all uses of the target word in all of its grammatical forms from the given corpus. Optionally, N uses may be sampled if there are more than that in order to reduce the following computations.
For intrinsic evaluation, instead of retrieving uses, golden uses may be taken, i.e., the uses which were shown to the annotators during dataset construction. This eliminates possible disagreements due to some rare senses of the target word sampled during the dataset annotation procedure but not sampled by the model during evaluation, or vice versa.
Build contextualized embeddings
(uses → embeddings). We employ BERT or XLM-R masked language model to obtain the contextualized representations of the given target word in the given text fragment (Devlin et al., 2019; Conneau et al., 2019). [^11] In principle, the benchmark supports all model checkpoints from huggingface. The simplest and most popular option is taking the outputs of the last Transformer layer (before the MLM head) on the positions of the target word and average those outputs (mean pooling). We can apply other subword poolings, max pooling or first pooling (take the outputs from the first subword). A more general implementation combines the outputs from several layers. Again, the simplest option is just averaging them. The experiments with BERT without fine-tuning presented in (Devlin et al., 2019) for the NER task suggest that it may be better to concatenate the outputs of the last four layers instead of averaging them. Thus, the layer aggregation function is a hyperparameter and we select its value among averaging and concatenation. Additionally, the benchmark supports XL-LEXEME, a WiC-fine-tuned version of XLM-R (Cassotti et al., 2023) and the similar XL-DURel which is instead fine-tuned for ordinal WiC (Yadav and Schlechtweg, 2025).
Generate pairs
(uses → pairs of uses). Pairs of uses of the specified type are generated. For the COMPARE type, each pair contains a use from the old corpus and a use from the new corpus. This type is ideal for the COMPARE task in which one needs to estimate the average similarity between old and new examples. If type is ALL, then all possible pairs are generated. This is useful to provide more information for the following clustering step.
For intrinsic evaluation, the golden pairs, i.e., those pairs that were annotated by humans during the dataset construction process, may be directly fed to the following components.
WiC
(pairs of uses → pairwise scores). We consider two different approaches to the Word-in-Context task. The first approach builds the contextualized embeddings for all uses with the aforementioned component, it inherits all the corresponding hyperparameters. Then it employs one of the distance functions to calculate distances between the contextualized embeddings of two uses in each pair. The euclidean, manhattan and cosine distances are currently supported.
The second approach employs a binary classifier implemented as a neural network that jointly processes a pair of word usages and returns the probability that the meaning is the same. We treat this probability as the similarity between word occurrences. We experimented with several such models that were part of the SOTA or near-SOTA solutions proposed by the DeepMistake team in the RuShiftEval and LSCDiscovery shared tasks (Arefyev et al., 2021a; Homskiy and Arefyev, 2022) or showed good performance on recent ordinal WiC shared task data (Yadav and Schlechtweg, 2025; Schlechtweg et al., 2025).
We further provide the possibility of discretizing graded WiC predictions at specified thresholds. By default, the thresholds from Yadav and Schlechtweg (2025) are used.
Clustering
(pairwise scores → clusters or contextualized embeddings → clusters). The goal of this step is to discover all senses of the target word occurring in the old, or the new corpus, or both of them, i.e., solve the WSI task for all uses retrieved from both corpora. Those clustering algorithms that can accept a matrix of similarities or distances between objects instead of object vectors can be applied to both types of inputs to the clustering component (pairwise scores, embeddings). However, some clustering algorithms require raw object vectors and can be applied to the contextualized embeddings, but not to the pairwise scores. For instance, K-Means calculates cluster centroids and needs object vectors to do that. The benchmark currently supports correlation clustering (Bansal et al., 2004) operating on pairwise scores. We chose this algorithm as most gold annotations were clustered with this approach (e.g. Schlechtweg et al., 2020).
Cluster measures
(clusters → LSCD predictions). To solve the binary LSCD tasks, after WSI we search for those clusters that contain only new or only old examples. Those clusters are viewed as novel or lost senses correspondingly, and the binary labels are predicted accordingly. Alternatively, we search for the clusters with at least M new examples and at most K old examples or vice versa in order to align with the annotation procedure of some LSCD datasets. To solve the JSD task, for old and new uses separately we estimate the probability distribution over senses of the target word as the proportions of its uses ended in each cluster. Then JSD between two probability distributions is calculated. Finally, to solve the COMPARE task we average the indicators that two uses ended in the same cluster for all pairs of uses of the COMPARE type.
Figure 4: Result overview on WiC.
Aggregate measures
(pairwise scores → LSCD predictions or contextualized embeddings → LSCD predictions). These skip the clustering step by aggregating WiC predictions directly. The most successful of these is Average Pairwise Distance (APD), which simply averages the distances or similarities between word usages from different time periods (COMPARE pairs) returned by some WiC model (Kutuzov and Giulianelli, 2020). This follows the calculation of the gold COMPARE scores, which is the average of human pairwise annotations. APD is sensitive to the polysemy or variation of the target word, which can lead to wrong predictions. For this, measures normalizing APD by a polysemy term were proposed, such as DiaSense (Beck, 2020). COS, in contrast, averages the contextualized embeddings of all old and all new uses separately producing two aggregated embeddings of the target word for each of two time periods, then calculate the distance between those two embeddings. It was previously under the names of PRT (Kutuzov and Giulianelli, 2020) and COS (Laicher et al., 2021). This distance is used as the prediction for both graded tasks.
7.  Experiments
We now use the benchmark the perform a number of experiments. We focus on Graded Change and COMPARE detection as these are the most widely approached LSCD tasks (Kutuzov and Pivovarova, 2021; Periti and Tahmasebi, 2024). Note that not all datasets provide both evaluation scores (see Table 1). All experiments are performed on the CoMeDi test split described in Section 1. We cannot test all models and configurations, thus we focus on SOTA models using aggregate change measures and design experiments to answer open research questions.
Which WiC model and which aggregate measure gives SOTA performance?
According to recent studies (Zamora-Reina et al., 2022; Periti and Tahmasebi, 2024; Zamora-Reina et al., 2025), DeepMistake and XL-LEXEME combined with APD compete for the SOTA on Graded Change and COMPARE detection. Previous studies have not directly compared these models though, or only on very limited data (Zamora-Reina et al., 2025). Hence, we perform a direct comparison guaranteeing a fair evaluation setup. We also include the recently published XL-DURel (Yadav and Schlechtweg, 2025), which has been shown to improve upon both models on ordinal WiC. In addition to the canonical datasets, we include a number of recently published datasets which have never be used for model evaluation: DWUG DE/EN/SV V3.0.0, DWUG DE/EN/SV resampled and DWUG IT. This is the most thorough model comparison done so far in the field of LSCD.
Figure 5: Result overview on dataset versions for bi-encoder models.
Figure 3 shows the results across the three SOTA WiC models, each combined with the common aggregate change measures APD, COS and DiaSense (see Section 6). Both, for Graded Change and COMPARE, clearly APD dominates which confirms previous results. However, there are slight advantages for the recently published XL-DURel model suggesting a new SOTA.
Does WiC prediction discretization improve results?
All above-reported measures based on aggregation of graded WiC predictions (either cosine similarity or same-sense probability). However, all WUG datasets are annotated on an ordinal (discrete) scale (see Section 4). Exactly predicting these ordinal values can be done by thresholding the graded predictions (Choppa et al., 2025). We hypothesize that discretizing WiC predictions to resemble human annotations helps for LSCD as ordinal judgments were used for ground-truth construction. Hence, in Figure 3 we report two models with APD and thresholding (XLD/XLM-TH-APD). Threshold parameters were taken from Yadav and Schlechtweg (2025). As we see, for XL-DURel, thresholding slightly helps, further pushing the SOTA, while for XL-LEXEME it has no effect or slightly hurts.
Which model gives SOTA on diachronic WiC? Does WiC determine LSCD?
A recent shared task has compared DeepMistake and XL-LEXEME for ordinal WiC showing a slight advantage for DeepMistake (Schlechtweg et al., 2025). [^12] Note that these models do not exactly correspond to the checkpoints we used. XL-DURel has further improved upon both models. Performance on the WiC task is a strongly influential factor for LSCD (Arefyev et al., 2021b). We compare model performance on these two levels to gain an understanding how strongly WiC determines LSCD performance. Figure 4 compares WiC models on the ordinal WiC task. The MCL checkpoint of DeepMistake does dominate XL-LEXEME confirming previous results, but both are outperformed by XL-DURel. Now compare this to the APD columns in Figure 3. Overall, WiC performance determines LSCD performance. Consider e.g. the dominance of XL-DURel on DWUG EN with graded change. However, there are notable exceptions such as DWUG SV where XL-DURel dominates on WiC but not LSCD. This shows that purely improving WiC ranking does not guarantee better LSCD performance. Possibly, the score distribution plays an important role when averaging values for aggregate measures.
Are model performances reproducible with more reliable data? What is the performance development on incrementally annotated datasets?
Many datasets have been annotated in incremental rounds of annotation (Schlechtweg et al., 2021, 2024a). Later rounds are supposed to yield higher data quality as graphs are more richly annotated. Schlechtweg et al. (2024a) suggest that previous model comparisons done on older dataset versions/less rounds should be repeated with the more reliable data (last round). We are the first to investigate model performance with the latest dataset versions. This will also allow us to investigate the impact of annotation rounds on performance. Figure 5 shows the performance of a selection of models on consecutive versions. For the top models, we can see that performance tends to increase with later versions, which is expected as quality should increase. However, there are cases where model performance strongly drops for newer versions, such as XLM-APD on DWUG EN. Further, the relative performance of models can strongly change depending on version, e.g. XLM-APD outperforms XLD-APD on DWUG EN for an older version while it is the other way around for the newer, more reliable version. This shows the risks of using unreliable ground-truth data and illustrates the need for the creation of benchmarks such as ours and continuous model reevaluation on additional and improved datasets.
Figure 6: Result selection on preprocessing. lemma = lemmatization, norm = normalization, raw = no preprocessing, token = tokenization, toklem = tokenization with target word substituted by lemma.
What is the impact of spelling variation on performance?
Laicher (2020) show that historical spelling variations can have a strong influence on BERT-based model performance. However, it is not clear how much this influences SOTA models. Hence, we test the influence of spelling normalization and lemmatization on two German datasets containing historical spelling variants in Figure 6. We can see that applying no preprocessing (raw) gives top performance for all models, with one exception (MCLen with toklem). This suggests that current base embedders are quite robust against spelling variation and do not need additional preprocessing. [^13] Note that ‘raw’ on these datasets already incorporates minimal preprocessing substituting a small number of historical characters with modern ones.
8.  Conclusion